Convert the following vectors;
75 m/s on a bearing of 295° to Cartesian form
[-3, -8] to direction/magnitude form
1
Expert's answer
2020-03-02T09:22:59-0500

x = r × cos( θ ) = 75 *  cos(295°) = 31.7 m/s

y = r × sin( θ ) = 75 * sin(295°) = -68.0 m/s

Answer: [31.7, -68.0]


r = √ ( x2 + y) = √ ( (-3)2 + (-8)2 ) = 8.54

θ = tan-1 ( y / x ) = tan-1 ( -8 / -3 ) = 69.4°

Answer: θ = 69.4°, r = 8.54